### v2

# Domain Adaptation from Generated Multi-Weather Images for Unsupervised Maritime Object Classification

Abstract.
The classification and recognition of maritime objects are crucial for enhancing maritime safety, monitoring, and intelligent sea environment prediction. However, existing unsupervised methods for maritime object classification often struggle with the long-tail data distributions in both object categories and weather conditions. In this paper, we construct a dataset named AIMO produced by large-scale generative models with diverse weather conditions and balanced object categories, and collect a dataset named RMO with real-world images where long-tail issue exists. We propose a novel unsupervised domain adaptation approach that leverages AIMO (source domain) to address the problem of limited labeled data, unbalanced distribution and domain shift in RMO (target domain). Specifically,we enhance the generalization of source features with the Vision-Language Models (i.e., CLIP) by Self-Knowledge Distillation. And we propose a difficulty score for curriculum learning to optimize training process by gradually introducing the generated data. Experimental results shows that the proposed method significantly improves the classification accuracy, particularly for samples within rare object categories and weather conditions. Datasets and codes will be publicly available.
Domain Adaptation, Maritime Object Classification
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 33rd ACM International Conference on Multimedia; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Object recognition
1. Introduction
The classification and recognition of maritime objects is essential for intelligent prediction of the sea environment, with wide applications in maritime traffic supervision, ship rescue and unmanned aerial vehicles (UAVs) cruising. Unsupervised maritime object classification based on transfer learning plays a pivotal role in enhancing maritime safety and monitoring by reducing the need for costly labeled data. In recent years, with the development of deep learning technology, image classification technology based on convolutional neural networks (CNN) and Vision Transformer (ViT) has been gradually applied to maritime object classification, achieving rapid development and breakthroughs.
Unsupervised maritime object classification presents significant challenges in handling the high similarity between ships and complex maritime environments. Large-scale annotated datasets are not common because of the difficulty in obtaining maritime objects and the high cost of manual labeling (Wang et al., 2020). Salem et al. (Salem et al., 2022) used fine-tuning strategies based on a pre-trained model and achieved good classification results, which also reduced training samples by half. Zhao et al. (Zhao et al., 2020) proposed the double transfer method to solve the problem of sample scarcity. However, the long-tail distribution of object categories, where a few categories dominate and the majority are underrepresented, still poses a critical hurdle.
Figure 1. Examples and data statistics of AIMO and RMO. (a): Examples of AIMO with multiple weather and illumination conditions. (b): Examples of RMO with multiple weather and illumination conditions. (c): The number of AIMO and RMO with multiple weather and illumination conditions. (d): The number of AIMO with different categories. (e): The number of RMO with different categories.
The performance of maritime object recognition is also complicated by adverse weather conditions, such as fog, rain, and low visibility. DTDNet (Liu and Zhou, 2022) proposed an effective dehazing method to guarantee reliable ship detection under foggy conditions. IRDCLNet (Sun et al., 2022b) aimed to improve the ship instance segmentation performance on proposed Foggy ShipInsseg dataset. However, the methods trained on specific weather phenomena usually fail to generalize across various maritime conditions. FREGNet (Tian et al., 2024) concentrated on reducing the impact of adverse weather and uneven illumination on supervised object recognition. Raza et al. (Raza et al., 2022) designed SimuShips dataset consisting of maritime objects under multiple weather and time conditions, but this work lacks specific categories for objects, which is only used for obstacle object detection. In summary, unsupervised maritime object classification under various weather conditions has not been well addressed yet, and new approaches need to be explored.
Due to the shortage of a large label-rich and balanced maritime object data under various weather and illumination conditions, we create AI-generated dataset for Maritime Object classification (AIMO), to compensate for the scarcity of weather conditions in real datasets and address the imbalance in the number of maritime objects across different categories. As shown in Fig. 1, AIMO is generated by Stable Diffusion, and naturally has a rich set of labels for categories and labels for multiple weather and illumination conditions. We also construct a real image dataset RMO (Real-world Maritime Object benchmark) by collecting from existing datasets and website. Even though advanced diffusion models generate images that are visually indistinguishable from real images, there are still significant differences between their distributions (You et al., 2024). Therfore, we propose a domain adaptation framework in which the generated data serves as the source domain, and the unlabeled real-world data acts as the target domain, facilitating improved classification accuracy under diverse weather conditions in real-world scenarios.
Considering that the generated data may suffer from insufficient diversity and poor generalization, we first enhance the feature representation of the source domain using CLIP (Radford et al., 2021). Specifically, we design text prompts by combining maritime object categories and weather conditions, and train an image feature extractor to be aligned with the image and text feature spaces of CLIP. Facing the unreliability of the generated source domain and the unlabeled target domain, inspired by Sun et al. (Sun et al., 2022a), we introduce adaptive perturbations to the features to prevent overfitting during training. Furthermore, we employ a dynamic curriculum learning strategy. We comprehensively consider AIMO’s weather and category information, and gradually select samples for training to make model easy to converge.
The contributions of this paper are summarized as follows:
  • We propose an effective approach to address the challenge of unsupervised maritime object recognition under various weather conditions, effectively handling adverse weather and mitigating the long-tail problem in maritime object classification.
  • We design an innovative domain adaptation framework for aligning the generated and real data, which overcomes the challenges caused by generated data with the proposed feature generalization enhancement module and curriculum learning strategy.
  • We construct two datasets, AIMO and RMO, which will contribute to advancing researches in related field.
2. Related Work
2.1. AI-generated Datasets
Currently, one of the most critical challenges in the advancement of Artificial Intelligence (AI) is the lack of high-quality data. As a result, there has been considerable attention on image generation technology, although the ongoing debate about the reliability of generated data persists.
Researcher at MIT discovered that, for the person capture task, models trained on generated data (videos with fewer background objects) outperform those trained on real data (Zewe, 2022). He et al. (He et al., 2023) identified that larger data volumes and greater data diversity are crucial for achieving better pre-training outcomes with generated data. Furthermore, from the perspectives of model architecture and pre-training methodologies, Vision Transformer (ViT)-based models tend to be more effective for pre-training on generated data than convolutional neural networks (CNNs). Additionally, self-supervised learning methods have been shown to be more suitable for pre-training with generated data than traditional supervised methods (He et al., 2023).
However, recent studies indicate that the use of AI-generated data may lead to model collapse (Shumailov et al., 2024). It is probably due to the generated data not being able to adequately represent the diversity and distribution of real data, which causes the model to gradually overfit these distorted data during training, ultimately compromising its ability to generalize, and leading to model collapse in the worst case. Model collapse not only diminishes accuracy, but also fosters self-reinforcing biases within the model, which further undermines the reliability of models. In light of this challenge, our research aims to utilize AI-generated data gradually from easy to difficult samples. We leverage the generated dataset, AIMO, for an unsupervised framework based on ViT in maritime object classification. Also, we utilized an Adaptive Adjustment Mechanism to address the issues of data overfitting. To the best of our knowledge, we are the first to overcome the long-tail issue in aspects of object categories and weather conditions for unsupervised maritime object classification by leveraging AI-generated data.
2.2. Unsupervised Domain Adaptation
Unsupervised domain adaptation (UDA) seeks to align the source and target domains by learning a feature representation that is invariant across domains. A series of method for align domains is to minimize the distribution divergence between source and target domains with discrepancy measures, i.e., maximum mean discrepancy (MMD) (Kang et al., 2019), central moment discrepancy (CMD) (Zellinger et al., 2017), correlation alignment (CORAL) (Sun et al., 2016) and maximum density divergence (MDD) (Li et al., 2020). Another kind of methods is motivated by adversarial learning, playing a two-player min-max game to learn domain-invariant representations, i.e. DANN (Ganin and Lempitsky, 2015), which introduced a domain discriminator to generate domain-invariant features. Moreover, self-training is used for domain adaptation to generate pseudo-label for the target domain and apply it in the training process (Mei et al., 2020).
With the rise of powerful Vision Transformer (ViT) for global feature extraction and modeling, an increasing number of research efforts are focusing on the application of transformers in UDA. ViT-base and ViT-small are from a pioneering work (Dosovitskiy et al., 2021) applying ViT to the task of image classification. Many works have been proposed on this basis. Swin (Liu et al., 2021), (Liu et al., 2022) focused on local self-attention computation and further proposed a large-scale pre-trained model serving as a generic backbone for computer vision. PMTrans (Zhu et al., 2023) used an intermediate domain to connect the source domain and the target domain, in order to bridge the domain gap, and conducted a min-max cross-entropy game on it. CDTrans (Xu et al., 2021) proposed cross–attention on source-target image pairs for direct feature alignment. TVT (Yang et al., 2023) proposed a multi-head self-attention module to obtain both transferable and discriminative features and combined it with adversarial adaptation.
Moreover, considering the strong generalization ability of Vision-Language Models (VLMs), a series of methods focus on the combination of UDA with VLMs. It is proved that VLMs are capable of handling multi-modal information (text and image) to further minimizes the domain gap. Many works have focused on fine-tuning the VLMs. Cheng et al. (Cheng et al., 2024) subsequently disentangled text prompts into domain-invariant and domain-specific descriptions. Li et al. (Li et al., 2024) designed a modality separation network to distinctly disentangle VLM’s image embeddings into language-associated and vision-associated components. However, directly fine-tuning these VLMs into downstream tasks may be computationally expensive, due to the large number of parameters (Lai et al., 2024). Prompt tuning is an efficient learning method with freezing both image and text encoders. Lai et al. (Lai et al., 2024) proposed a domain-aware pseudo-labeling scheme designed for VLMs to achieve domain disentanglement. Bai et al. (Bai et al., 2024) proposed a Prompt-based Distribution Alignment (PDA) method to integrate the domain knowledge into prompt learning. DAMP (Du et al., 2024) learned prompts mutually with a cross-attention module based on text embeddings and image embeddings extracted from VLM.
Different from the above methods, we utilize a self-knowledge distillation method to transfer the knowledge of VLMs in a flexible and stable way. We distill the knowledge of text embeddings and image embeddings extracted by frozen VLMs into feature extractor to enhance feature representation, which simultaneously align semantic space and visual space by carrying out bidirectional constraints.
Figure 2. An overview of the proposed method. We use a series of Vision Transformer Blocks as backbone of feature extraction and take labels (classes and weather conditions) and images from the source domain, as well as images from the target domain as inputs. The designed UDA framework for maritime object classification is an adversarial adaptation network, consisting of generalization enhancement for source features, adversarial adaptation from source to target features and classifier refinement with perturbed features from both domains.
3. Method
We aim to enhance the unsupervised maritime object classification in real-world scenarios by leveraging the generated data. The proposed adversarial UDA framework is shown in Fig. 2, which transfers the knowledge from labeled source domain (balanced generated data) to unlabeled target domain (unbalanced data in reality). The main modules contain generalization enhancement, adversarial adaptation and classifier refinement.
Specifically, the Generalization Enhancement aligns the features from Transformer Blocks with the text-encoder and image-encoder features of CLIP for source domain, which could be regarded as a Self-Knowledge Distillation process. The inputs of CLIP are the source domain data with rich information of category and multiple weather conditions. We measure the similarity between text embeddings, image embeddings and source domain features to constrain this process. Inspired by Sun et al. (Sun et al., 2022a), Classifier refinement improves the prediction accuracy by enhancing the robustness of feature representation. Random offsets are added into the input token sequence of a randomly chosen transformer block, which is performed on images from both source and target domains to acquire perturbed features. Then we employ the Kullback-Leibler (KL) divergence to quantify the discrepancy of the corresponding predicted class probabilities of original and perturbed features, and carry out random offset refinement. Furthermore, we use an Adaptive Adjustment Mechanism to carry out the training process step by step by scoring the source domain data.
3.1. Preliminary
We consider the image classification task in UDA, where a labeled source domain $\mathcal{D}_{s}={\{(\bm{x}^{s}_{i},\bm{y}^{s}_{i})\}}^{n_{s}}_{i}$ with $n_{s}$ examples and an unlabeled target domain $\mathcal{D}_{t}={\{\bm{x}^{t}_{i}\}}^{n_{t}}_{i}$ with $n_{t}$ examples are given. Note that the two domains share the same label space $\mathbb{R}^{K}$, where K is the number of classes.
The basic adversarial UDA framework is composed of three parts: feature extractor F, consisting of a Patch Embedding layer and a series of Transformer Blocks, classifier C, and discriminator D. Given an input $\bm{x}$ from either source domain or target domain, $\bm{S}=\textbf{{F}}(\bm{x})$ is referred to a token sequence encoded by F, which is passed to D directly. Our novel adversarial UDA framework contains a Generalization Enhancement module G for enhancing the source domain feature. Therefore, the source domain feature $\bm{S}_{\bm{x}^{s}}$ is passed to G, and then $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}}=\textbf{{G}}(\bm{S}_{\bm{x}^{s}})$ is passed to C. The C predict $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}}$ into class probabilities $\bm{p}=\textbf{{C}}(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}})\in\mathcal{R}^{K}$. Therefore, the classification loss on source domain data is computed as
(1) $\mathcal{L}_{cls}=-\mathbb{E}_{(\bm{x},\bm{y})\in\mathcal{D}_{s}}\sum_{k=1}^{K}\mathbb{1}_{k=y}\log{\bm{p}_{k}},$
which is actually a standard cross-entropy loss.
Meanwhile, D predicts $\bm{S}$ into domain logits $\bm{q}=\textbf{{D}}(\bm{s})\in\mathcal{R}^{2}$, which is actually a binary domain discrimination to learn domain-invariant feature. We define the domain adversarial loss $\mathcal{L_{D}}$ as
(2) $\mathcal{L}_{dom}=-\mathbb{E}_{\bm{x}\in\mathcal{D}_{s}}\log{\bm{q}_{s}}-\mathbb{E}_{\bm{x}\in\mathcal{D}_{t}}\log{\bm{q}_{t}},$
where $\bm{q}_{s}$ and $\bm{q}_{t}$ represent source and target domain, respectively. To achieve domain alignment, a domain-invariant feature encoder is needed to confuse the domain discriminator. Therefore, we adversarially train D to minimize $\mathcal{L}_{dom}$ and F to maximize $\mathcal{L}_{dom}$, which is implemented by reversing the gradients flowing from D to F.
The objective of UDA is formulated as
(3) $\min_{\textbf{{F}},\textbf{{C}}}\max_{\textbf{{D}}}\mathcal{L}=\mathcal{L}_{cls}-\mathcal{L}_{dom}+\alpha\mathcal{L}_{SKD}+\beta\mathcal{L}_{offset},$
where $\alpha$ and $\beta$ are trade-off parameters introduced in Sec 3.4. $\mathcal{L}_{SKD}$ represents the Self-Knowledge Distillation (SKD) loss introduced in Sec 3.2 and $\mathcal{L}_{offset}$ represents the random offset refinement loss on both source and target domain introduced in Sec 3.3.
3.2. Generalization Enhancement Based on CLIP
VLMs exhibit exceptional performance across various distributions, attributed to the vast diversity of distributions seen during their training (Radford et al., 2021), which leverage contrastive pre-training methods (Zhai et al., 2022) for learning the correlation between the embeddings of matching image-text pairs. For instance, CLIP (Radford et al., 2021) are trained jointly on 400 million image-text pairs, yielding remarkable generalization across different data distributions. Despite the considerable zero-shot generalization of VLMs in classification tasks, effectively adapting them to downstream tasks remains a great challenge. It is expected to leverage the knowledge of VLM in a portable way. Knowledge distillation not only avoids large-scale parameter tuning of VLMs, but also helps to realize lightweight reasoning. For UDA task, it is an effective strategy that enhances cross-domain representation across the open semantic and visual space of VLMs, in order to bridge the domain gap. Compared to a single label, VLMs have more descriptive information, which is beneficial of making full use of the AIMO’s rich label information as semantic consistency constraint. Therefore, we propose a Generalization Enhancement module to efficiently distill the abundant knowledge from CLIP into the adversarial UDA network. We distill the knowledge of text embeddings and image embeddings extracted by frozen VLMs into feature extractor of ViT to enhance feature representations, by carrying out a two-branch constraint to simultaneously align backbone features with semantic information and visual information from VLMs.
Specifically, we perform a Self-Knowledge Distillation on the source domain data based on CLIP. The feature invariance is constructed in the source domain by Self-Knowledge Distillation, which makes the model more robust to the distribution migration in the target domain. Firstly, We utilize the frozen text encoder and image encoder of VLMs to extract embedded features, which update the source domain image representations as a ”teacher signal”. And then the similarity between text embeddings, image embeddings and backbone features is calculated respectively, as constraints to supervise the training process, in order to realize a double alignment of semantic space and visual space. Given labels for category and weather conditions, as well as images from the source domain as inputs to CLIP, the source domain features of Transformer Blocks are aligned with the text-encoder and image-encoder features of CLIP in the Generalization Enhancement module. As proper text prompt allow for better control of invariance in text embeddings (Addepalli et al., 2024), we use ”A photo of {class} in {domain} to represent the semantic embedding. {domain} is referred to the weather and illumination condition of the source image. We denote the CLIP’s text embedding for ”A photo of {class} in {domain} and image embedding of the source domain input $(\bm{x}^{s}_{i},\bm{y}^{s}_{i})$ as $\bm{T}_{{y}^{s}_{i}}$ and $\bm{I}_{\bm{x}^{s}_{i}}$. The generalization enhanced feature corresponding to the input $\bm{x}^{s}_{i}$ are denoted as $\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}}$. The following Self-Knowledge Distillation loss is used to refine the aligned features:
(4) $\mathcal{L}_{SKD}=-\mathbb{E}_{(\bm{x},\bm{y})\in\mathcal{D}_{s}}\sum_{i=1}^{n_{s}}\{\cos(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}},\bm{T}_{{y}^{s}_{i}})+\cos(\bm{S}^{\textbf{{G}}}_{\bm{x}^{s}_{i}},\bm{I}_{\bm{x}^{s}_{i}})\}.$
where $\cos(\cdot,\cdot)$ denotes the cosine similarity between two vectors, and the objective of $\mathcal{L}_{SKD}$ is to distill the comprehensive image features and semantic representations learned by the image encoder and text encoder of the CLIP.
3.3. Classifier Refinement
We regularize the latent feature spaces of the transformer backbone through classifier refinement using perturbed source and target domain features to reduce overfitting . Incorporating a random offset into the data is a prevalent data augmentation technique that is widely employed to alleviate overfitting. It may be better to perform data augmentation operations at hidden layers rather than directly manipulating the input images (Verma et al., 2019). The optimal layer for introducing perturbations differs from various tasks. Perturbing relatively deep layers tends to yield superior performance but increases the risk of model collapse, and randomly choosing one transformer block proves to be more robust compared with perturbing any individual layer from them (Sun et al., 2022a). Therefore, we select a transformer block at random, and then add a random offset to the input token sequence of the chosen block.
We perform regularization on both the source and target domains. Given an image $\bm{x}_{i}$ from either source domain or target domain, $\bm{S}^{l}_{\bm{x}_{i}}$ represents its input token sequence of the l-th transformer block. $\bm{S}^{l}_{\bm{x}_{i}}$ can be regarded as a latent representation of $\bm{x}_{i}$ within a hidden space. And then we utilize the token sequence $\bm{S}^{l}_{\bm{x}_{j}}$ of another randomly chosen image $\bm{x}_{j}$ to add an offset. The perturbed token sequence of $\tilde{\bm{S}}^{l}_{\bm{x}_{i}}$ is obtained as
(5) $\tilde{\bm{S}}^{l}_{\bm{x}_{i}}=\bm{S}^{l}_{\bm{x}_{i}}+\gamma(\bm{S}^{l}_{\bm{x}_{j}}-\bm{S}^{l}_{\bm{x}_{i}}),$
where $\gamma$ is a trade-off parameter introduced in Sec 3.4. Note that gradients do not back-propagate through the offset, while they pass through $\bm{S}^{l}_{\bm{x}}$.
We use KL divergence to measure the distance of the predicted probability vectors $\bm{p}$ and $\bm{\tilde{p}}$, corresponding to $\bm{S}^{l}_{\bm{x}}$ and $\tilde{\bm{S}}^{l}_{\bm{x}}$, respectively. The following equation shows the KL divergence between the perturbed predicted probability and the original predicted probability.
(6) $\bm{D}_{\text{KL}}(\bm{p}||\bm{\tilde{p}})=\sum_{i=1}\bm{p}(i)\log\frac{\bm{p}(i)}{\bm{\tilde{p}}(i)}.$
$\bm{D}_{\text{KL}}(\bm{\tilde{p}}||\bm{p})$ is the calculated similarly. And it is more robust to combine the two KL divergence together in random offset loss (Sun et al., 2022a). To achieve reliable predictions, we define a Confidence Filter $\mathcal{F}$ as
(7) $\mathcal{F}[\mathcal{D};\bm{p}]=\{\bm{x}\in\mathcal{D}|\max(\bm{p})>\kappa\},$
where $\kappa$ is predefined threshold.
The random offset loss is defined as an average on every training batch $\mathcal{B}\subseteq\mathcal{D}_{\bm{s}}\cup\mathcal{D}_{\bm{t}}$, containing images from source domain and target domain.
(8) $\begin{split}\mathcal{L}_{offset}=\mathbb{E}_{\mathcal{B}\subseteq\mathcal{D}_{\bm{s}}\cup\mathcal{D}_{\bm{t}}}\{\omega\mathbb{E}_{\bm{x}\in\mathcal{F}[\mathcal{B};\bm{p}]}\bm{D}_{\text{KL}}(\bm{p}||\bm{\tilde{p}})\\ +(1-\omega)\mathbb{E}_{\bm{x}\in\mathcal{F}[\mathcal{B};\bm{\tilde{p}}]}\bm{D}_{\text{KL}}(\bm{\tilde{p}}||\bm{p})\},\end{split}$
where $\omega$ is a random variable drawn from a Bernoulli distribution $\mathcal{B}(0.5)$ (Sun et al., 2022a). Note that the loss gradient is back-propagated through both the perturbed and the original predicted probabilities.
In addition to the random perturbation, the multimodal information is introduced only for the unperturbed source domain images in Sec 3.2 and the Dropout layer within the classifier head operates in a stochastic manner for both branches, which introduce additional source of discrepancy for the random offset refinement loss.
3.4. Adaptive Adjustment Mechanism
Domain adaptation using synthetic data remains a challenging task, especially for long-tailed distributed maritime objects across diverse weather conditions and multiple categories, where substantial variations in sample difficulty exist. Accordingly, we propose an Adaptive Adjustment Mechanism based on curriculum learning and adversarial training, in order to avoid model falling into local optimal, reduce data overfitting and avoid possible model collapse. By imitating the ”step by step” learning process of human, the curriculum learning gradually guides the adversarial training to realize the alignment of source domain and target domain. We make models quickly establish basic representations by firstly feeding easier samples, and then gradually introducing the harder ones. Traditional curriculum learning relied on prior knowledge (usually using fixed values with experience) to evaluate sample difficulty, which leads to lower effectiveness in the later period of training (Wang et al., 2021). Recent method (Jiang et al., 2025) tried using losses as dynamic difficulty evaluation. In this paper, we propose the difficulty score for our task as a combination of prior score and dynamic loss (Eq. LABEL:comprehensive_score) to select source domain samples (i.e., the generated data) for training.
Firstly, we design a prior score as the Image Quality Assessment (IQA) score weighted by weather condition. Specifically, we utilize CLIP IQA (Chen and Mo, 2022) for the prior IQA analysis of the source domain images. According to different weather conditions, the pre-defined weights are:{sunny: 5, sunset & night: 4.5, cloudy: 4, foggy:3.5, rainstorm: 3}. Subsequently, higher image quality and better weather condition contribute to a higher prior score. Secondly, the rest part of difficulty score is based on the standard cross-entropy loss (Eq. 1), where more certain category prediction leads to higher dynamic score. The difficulty score is formulated as follows:
(9) ${difficulty}\_{score}=(1-\phi){prior}\_{score}+\phi(1-\mathcal{L}_{cls}),$
where the initial value of the adaptive weight $\phi$ is 0 and gradually increases to 1 at the end of the training in a linear function.
Therefore, the selected source domain data for the t-th epoch is:
(10) ${A}_{t}=\operatorname{argmax}_{\hat{A}:|\hat{A}|\geq\lambda(t)\cdot|A|}{difficulty}\_{score},$
where $A$ denotes the source domain data, the $|A|$ and $|\hat{A}|$ represent the size of $A$ and $\hat{A}$. $\lambda(t)$ is used to schedule the proportion of source training samples at the t-th epoch, which is defined as:
(11) $\lambda(t)=\lambda_{0}+(1-\lambda_{0})(1-e^{-kt/T}),$
where $\lambda_{0}$ is the initial proportion of source training samples, which is set to 0.5, and $T$ is the total number of epochs. Note that $k$ is set to 2 for the amount of data increasing rapidly in the early stage of training and slowly in the late stage of training, in order to avoid making it difficult to generalize the model to difficult samples due to long-term training of high-score samples.
Furthermore, in the classifier refinement strategy, the selection of appropriate values for the perturbation scalar $\gamma$ and the random offset refinement loss weight $\beta$ is crucial. Excessively large perturbations result in a collapse of the predicted class distribution, whereas small perturbations may fail to fully leverage their potential benefits. Given that the target domain is entirely unlabeled and domain adaptation tasks vary significantly even within the same dataset, it is desirable to adjust these parameters adaptively. At the initial stage of training, the model is not stable, and too much perturbation may lead to convergence difficulties. With the increase of the number of iterations, the proportion of disturbance should be appropriately increased to reduce overfitting. Inspired by SSRT (Sun et al., 2022a), we plan to introduce random perturbations dynamically, by setting an an adaptive scalar.
Specifically, the adaptive scalar $\mu\in[0,1]$ is defined as
(12) $\mu(n)=\begin{cases}\sin\left(\frac{\pi}{2N}{n}\right)&\text{if }n<N\\ 1.0&\text{otherwise}\end{cases},$
where n points to the current training iteration, and N is a fixed period, setting to 1000. Therefore, r gradually goes up to 1.0 after N steps. The adaptive scalar $\mu$ is adopted to modulate $\gamma$ and $\beta$, i.e., $\gamma_{\mu}=\mu\gamma$ and $\beta_{\mu}=\mu\beta$, which builds a bridge between the adaptive adjustment and random offset refinement, making our proposed UDA framework more resilient to avoid data overfitting.
4. Experiments
4.1. Dataset Construction
To overcome the issues of single weather condition and long-tail distribution of maritime objects in existing datasets, we construct an AI-generated dataset, AIMO, and a real-world dataset, RMO with labels of diverse weather conditions and object categories.
AIMO consists of 66,626 images with category labels and multiple weather and illumination conditions. Taking advantage of the strong generation capabilities of Stable Diffusion-WebUI [^1] https://github.com/AUTOMATIC1111/stable-diffusion-webui.git, we create most of the maritime objects using simple text prompts. In order to make the generated data closer to the real data, we utilize techniques such as LoRA, image embeddings (Gal et al., 2022) and hypernetworks (Ha et al., 2016) to assist image generation. For instance, the generated results of warship series, e.g., aircraft carriers, are commonly not satisfactory, so we utilize a LoRA model  [^2] https://www.seaart.ai/models/detail/f8078759a92ae10610a92f8683f893da for better generation quality. The use ratio of LoRA is manually adjusted, according to the generation quality of different categories. In addition, for some categories that are beyond the coverage of the generated model, e.g., barges, tugs and so on, we select a small amount of real maritime objects images for specific categories to make image embeddings by extracting features. The produced image embeddings serve as a prior pseudo-prompt in the generation process, in order to enhance the authenticity and diversity of the generated samples. Considering the complexity of the sea surface environment, we train a hypernetwork on WSODD (Zhou et al., 2021) to achieve sea surface environment style patches, which make the generated data cover a wider range of weather and illumination conditions. As a small neural network, hypernetworks dynamically adjust the weight parameters of the main model (i.e., Stable Diffusion).
RMO consists of 32,418 images collected from multiple maritime object datasets and websites. Specifically, we exclude images that lack the object, as well as those depicting the object in sketch form. And for images containing multiple objects, we crop out each object separately.
The proposed AIMO involves five kinds of weather and time conditions, namely sunny, cloudy, foggy, rainstorm, and sunset & night. AIMO shares the same label space as RMO, covering 15 categories: aircraft carrier, barge, cruise ship, destoryer, ferry boat, fishing boat, freight ship, inflatable boat, lighthouse, maritime buoy, motorboat, pleasure boat, sailboat, submarine and tug. In order to better verify the effect of AIMO in real scenarios with different weather and illumination conditions, we manually divided the RMO into the above conditions. Note that the annotations are just used for validation where the labels are not accessible during training.
For the UDA task, AIMO and RMO are used as source domain and target domain, respectively. We use the average classification accuracy on RMO as the evaluation metric.
4.2. Training Details
We initialize ViT-B/16 (Dosovitskiy et al., 2021) with ImageNet (Russakovsky et al., 2015). We use the SGD optimizer (Le et al., 2011) with a constant learning rate of 0.002 for all our experiments and we set batch size to 16. Our backbone network contains 12 layers of ViT blocks, and we choose blocks 0, 4, and 8 as alternate layers. For features of each batch, a perturbation layer is randomly selected from the alternate layers, and the offsets are respectively added for the source domain and the target domain features in this layer. We set the hyper parameters $\alpha=0.3$, $\beta=1$, $\gamma=0.2$, and $\kappa=0.4$, according to the discussion in Sec 4.6.
Figure 3. Validation experiments with our proposed method and SCAN. (a): Comparison of accuracy on RMO with multiple weather & illumination conditions. (b): Comparison of accuracy on RMO with different categories.
4.3. Effectiveness of AIMO
In this section, we evaluate the image quality of AIMO, and design experiments to verify AIMO has the ability to improve the accuracy of maritime object classification in complex environments, and solve the classifier imbalance caused by the long-tail effect.
Firstly, we selected three no-reference image quality assessment metrics (Chen and Mo, 2022), ENTROPY, NIQE, and CLIPIQA to evaluate the quality of AIMO . ENTROPY is used to measure the complexity and information of images. NIQE is used to assess the natural quality of an image. CLIPIQA is used to evaluate image quality by using CLIP. Higher values of these metrics indicate better image quality. Also, we calculated the FID (Chen and Mo, 2022) metric between AIMO and RMO, which characterizes the difference in feature distribution. The results are shown in Table 1. ENTROPY over 7 and CLIPIQA close to 0.8 indicate that AIMO images have rich details, which contain valuable information to transfer. However, the lower NIQE and higher FID indicate a clear difference between AIMO and natural maritime object images, which requires effective domain adaptation.
In Fig.3, we compare our proposed method with SCAN (Van Gansbeke et al., 2020), an unsupervised clustering method that works only on unlabeled RMO. Real-world maritime object images are rarely taken under bad weather or uneven illumination conditions, and have obvious long-tail issues. Therefore, only using real-world data for training easily leads to unbalanced classification results.
Comparison of accuracy on RMO with multiple weather & illumination conditions. As shown in Fig. 3(a), the UDA method we proposed using AIMO has achieved better performance under various weather and illumination conditions. Particularly in rainstorm weather, our method has witnessed a significant performance boost. Moreover, in foggy weather with insufficient lighting and potential object occlusion, the classification accuracy of our method is over 18 % higher than that of the unsupervised classification method solely on RMO. Additionally, the introduction of AIMO has also improved the classification accuracy of maritime objects in other scenarios with relatively abundant lighting. The results indicates that the application of generated data containing richer environmental information enhances the classification performance under various situations, which prove that the effectiveness of AIMO with multiple weather and time conditions.
Table 1. Image quality assessment on AIMO
Dataset ENTROPY NIQE FID CLIPIQA
AIMO 7.325 4.874 34.011 0.784
Table 2. Comparison with other UDA method
Method Acc %
ViT-small (Dosovitskiy et al., 2021) 56.456
ViT-base (Dosovitskiy et al., 2021) 60.693
Swin-B (Liu et al., 2021) 58.761
CDTrans (Xu et al., 2021) 59.183
PMTrans (Zhu et al., 2023) 65.463
TVT (Yang et al., 2023) 62.260
PDA (Bai et al., 2024) 65.891
Ours 73.315
Comparison of accuracy on RMO with different categories. In Fig. 3(b), the UDA method we proposed using AIMO has achieved better performance across all categories. In the categories of pleasure yacht, destroyer, inflatable boat, lighthouse and maritime buoy, our method significantly improved the classification accuracy. We obverse that especially in the categories of inflatable boat, lighthouse and maritime buoy, the classification accuracy of our method was notably different from that of the unsupervised clustering method, because these three categories account for a relatively small proportion in RMO, as shown in Fig. 1(e). It is proved that by using the category balanced AIMO for transfer learning, we effectively improve the classification accuracy of categories with rare samples and alleviate the long-tail effect to a certain extent.
4.4. Effectiveness of Domain Adaptation
We compare our proposed method with other UDA methods, e.g., ViT-small (Dosovitskiy et al., 2021), ViT-base (Dosovitskiy et al., 2021), Swin-B (Liu et al., 2021), CDTrans (Xu et al., 2021), PMTrans (Zhu et al., 2023), TVT (Yang et al., 2023), and PDA (Bai et al., 2024) in Table 2.
Comparison with the basic networks. ViT-small (Dosovitskiy et al., 2021) and ViT-base (Dosovitskiy et al., 2021) have applied Transformer to image classification for the first time, demonstrating the potential of global attention in visual tasks. On this basis, Swin Transformer (Liu et al., 2021) takes into account global modeling and local concentration bias through windowing and hierarchical design. Compared with these basic networks, our proposed method gained of over 10 % on the classification accuracy, which attributes to our effective designs based on a series of modules in order to mitigate domain divergence.
Comparison with methods using pseudo labels. CDTrans (Xu et al., 2021) learned domain-invariant features through cross-domain attention and used target domain pseudo labels for self-supervised learning. Given the substantial differences in distribution between AIMO and RMO datasets, directly aligning image pairs from the source and target domains is not an effective approach. And the large domain gap makes the pseudo labels more unreliable, which directly affects the model performance. Therefore, CDTrans (Xu et al., 2021) exhibits suboptimal performance. Also, different from PMTrans (Zhu et al., 2023), which constructed intermediate domains to alleviate domain gap and also relied on the pseudo labels during training, our method utilizes VLMs to alleviate the domain-shift from source domain to target domain, achieving nearly 7.85 % gains on the accuracy.
Comparison with method using adversarial adaptation. TVT (Yang et al., 2023) designed a transferability adaption module based on adversarial adaptation. The learned transferabilities were injected into attention blocks, promoting ViT to focus on both domain-invariant and domain-specific features. Compared with TVT (Yang et al., 2023), which also utilizes adversarial training for UDA, the accuracy of our proposed method is 10 % higher, probably due to the use of VLMs (i.e., CLIP) for improving robustness of the source domain features.
Comparison with method using VLMs. PDA (Bai et al., 2024) performed a two-branch prompt tuning. Firstly, the zero-shot capability of CLIP is used to make pseudo labels of target domain, and further extracted the text embeddings and image embeddings of source domain and target domain. And then they used image embedding to guide the backbone network to learn the self-enhanced features and cross-domain features. Different from PDA (Bai et al., 2024), we achieved better classification results which benefits from injecting the semantic information into backbone network. Compared with PDA (Bai et al., 2024), which also utilizes VLMs, our proposed method performs more than 7.42 % on the classification accuracy.
Table 3. Ablation study for classification accuracy on RMO
Adversarial Training Input Offset Token Offset Generalization Enhancement Adaptive Adjustment Acc (%)
61.614
67.253
69.385
72.287
73.315
4.5. Ablation Study
For better understanding of the effect of each proposed module, we conduct an ablation study on RMO and show the result in Table 3.
We combine ViT-B/16 (Dosovitskiy et al., 2021) with adversarial training as the baseline and obtain a 61.612% on the classification accuracy of RMO. Adding offsets on input images directly makes a 5.639% increase compared to baseline in classification accuracy. Comparison of the second and third rows in Table 3, replacing the direct offset on images with adding random offset to the token sequence further improves the classification accuracy by 2.132%. It proves that injecting perturbations into the hidden space is a better choice. On this basis, the Generalization Enhancement Module is introduced to make the classification accuracy finally rise to 72.287%, which shows that it is meaningful to use CLIP for Self-Knowledge Distillation on the source domain. Finally, we employ curriculum learning to regularly select the source domain samples for training, which further improves the classification accuracy to 73.315 %. The result demonstrates the effectiveness of Adaptive Adjustment Mechanism. In summary, the ablation studies conducted on the RMO effectively validate the contributions of the proposed modules.
4.6. Parameter Sensitivity Discussion
In this part, we analyze the sensitivity of three key parameters in our framework: $\alpha$ and $\beta$ used in Eq. 3, and $\gamma$ in Eq. 5. The classification accuracy trends under varying parameters are visualized in Fig. 4. As shown in Fig. 4(a), the stable variation validates the robustness of the proposed Generalization Enhancement module. Fig. 4(b) reveals a critical dependency on $\gamma$, which underscores that the necessity of sufficient regularization that overly small $\gamma$ fails to mitigate overfitting, while larger values retain competitive accuracy. And the results in Fig. 4(c) imply that insufficient weight weaken its correction effect, while overemphasis may distort the primary learning objective. Therefore, the best choices for these parameters are $\alpha=0.3$, $\beta=1$, and $\gamma=0.2$.
Figure 4. Parameter sensitivity experiments.
5. Conclusion
In this work, we propose a label-rich and balanced AI-generated Benchmark for Maritime Object Classification (AIMO) with multiple weather and illumination conditions. And we construct a Real Maritime Object Benchmark (RMO) collected from a series of real maritime object datasets and the Google. We further introduce an UDA framework with adversarial training and curriculum learning based on Vision Transformer to address maritime object classification in the real-world scenarios. It leverages the generalization of VLM to refine source domain features and narrow domain gap. Experiments validate the effectiveness of the proposed benchmark and unsupervised domain adaptation framework.